# All loop soft photon theorems and higher spin currents on the celestial sphere

### v2

@@ Abstract
Soft factorization theorems can be reinterpreted as Ward identities for (asymptotic) symmetries of scattering amplitudes in asymptotically flat space-time. In this paper we study the symmetries implied by the all loop soft photon theorems when all external particles are massless. Loop level soft theorems are qualitatively different from the tree level soft theorems because loop level soft factors contain multi-particle sums. If we want to interpret them as Ward identities then we need to introduce additional fields which live on the celestial sphere but do not appear as asymptotic states in any scattering experiment. For example, if we want to interpret the one-loop exact $\mathcal{O}(\ln\omega)$ soft theorem for a positive helicity soft photon (with energy $\omega$) as a Ward identity then we need to introduce a pair of antiholomorphic currents on the celestial sphere which transform as a doublet under the $SL(2,\mathbb{R})_{R}$. We call them dipole currents because the corresponding charges measure the monopole and the dipole moment of an electrically charged particle on the celestial sphere. More generally, the soft photon theorem at $\mathcal{O}(\omega^{2j-1}(\ln\omega)^{2j})$ for every $j\in\frac{1}{2}\mathbb{Z}_{+}$ gives rise to $(2j+1)$ antiholomorphic currents which transform in the spin-$j$ representation of the $SL(2,\mathbb{R})_{R}$. These currents exist in the quantum theory because they follow from loop level soft theorems. We argue that under certain circumstances the (classical) algebra of the higher spin currents is the wedge subalgebra of the $w_{1+\infty}$.
@@ I Introduction
In this paper, we study the symmetry interpretation of the soft photon theorems that remain valid even at loop level. Consequently, the symmetries implied by these soft theorems are genuine symmetries of the holographic dual celestial CFT2, persisting even after quantum corrections are taken into account. It is well known that the leading soft photon theorem, derived in Gell-Mann:1954wra ; Low:1954kd ; Weinberg:1965nx , is universal and exact at tree level, receiving no loop corrections (up to a subtlety in pQCD with massless quarks as pointed out in Ma:2023gir ). However, since four-dimensional $S$-matrices are infrared divergent due to the long-range interactions mediated by massless particles, deriving finite subleading corrections has long remained a challenge Bern:2014oka ; He:2014bga . By interpreting soft factors as ratios of $S$-matrices with and without an external soft photon, and by systematizing the exponentiation of infrared divergences, it is possible to extract well-defined subleading soft theorems from the IR-finite parts of the ratios of $S$-matrices. This program was successfully carried out in Sahoo:2018lxl , with a proof of universality subsequently established in Krishna:2023fxg . The resulting subleading soft photon theorem appears at $\mathcal{O}(\ln\omega)$ in the soft photon energy $\omega$ and is one loop exact. The $\mathcal{O}(\ln\omega)$ soft factor consists of two distinct contributions, termed the “classical” and “quantum” parts in Sahoo:2018lxl . The classical contribution encodes information about the classical electromagnetic waveform in the low-frequency expansion, as demonstrated through independent classical analyses in Laddha:2018myi ; Saha:2019tub . By contrast, a clear physical interpretation of the quantum contribution remains elusive to date. An intriguing feature of the subleading soft photon theorem is that the classical contribution vanishes for massless charged particles, while the quantum contribution remains non-zero. This observation motivates a deeper investigation of the quantum part of subleading soft photon theorem in massless QED. In particular, we aim to understand it as a Ward identity in the framework of celestial CFT and to uncover the associated symmetry structure.
@@
From a four-dimensional bulk perspective, the classical part of the subleading soft photon theorem at $\mathcal{O}(\ln\omega)$ has already been understood as a Ward identity associated with a large gauge symmetry. [^1] The Celestial OPE of gluons and gravitons at one-loop were studied in Bhardwaj:2022anh ; Krishna:2023ukw . This symmetry acts on charged matter fields as a local enhancement of global phase rotations, referred to as “superphaserotations” in Campiglia:2019wxe ; AtulBhatkar:2019vcb ; Choi:2025mzg . In the present work we provide an interpretation of the quantum contribution to the $\mathcal{O}(\ln\omega)$ soft photon theorem in massless QED as a symmetry of the dual Celestial CFT2. Furthermore, we analyze higher subleading corrections to the soft photon theorem at orders $\omega^{n}(\ln\omega)^{n+1}$ as conjectured and explicitly derived up to sub-subleading order in Sahoo:2020ryf ; Karan:2025ndk , and argued to be $(n+1)$-loop exact. Analogously to the subleading soft photon theorem, we argue that at higher subleading orders only the quantum contributions survive in massless QED.
@@
Loop level soft theorems are qualitatively different from the tree level ones. It turns out that if we want to interpret them as Ward identities of symmetries in a Celestial CFT2 then we need to introduce additional fields on the celestial sphere which do not appear as asymptotic states in scattering experiments. In particular, they are not related to the soft particle in any obvious manner. Now since there are an infinite number of loop level soft photon theorems we get an infinite number of chiral currents in this way. For every $j\in\frac{1}{2}\mathbb{Z_{+}}$ the soft theorem at $\mathcal{O}(\omega^{2j-1}(\ln\omega)^{2j})$ gives rise to $(2j+1)$ chiral currents which transform in the finite dimensional spin-$j$ representation of the $SL(2,\mathbb{R})_{R}$. The charges corresponding to the $j=1/2$ currents measure the monopole and the dipole moment of an electrically charged particle on the celestial sphere and hence called a dipole current. The other spin-$j$ currents can be constructed by taking the (normal ordered) product of $2j$ dipole currents. Depending on the value of a real parameter $k$ which appears in the OPE of the dipole currents, the algebra of the higher spin currents can be Abelian ($k=0$) or the wedge subalgebra ($k<<1$) of the $w_{1+\infty}$.
@@
Appearance of higher spin currents is not new in the context of Celestial holography. Both soft gluons and gravitons generate infinite dimensional higher spin algebras known as the $S$-algebra and the wedge subalgebra of $w_{1+\infty}$ Strominger:2021mtt ; Adamo:2021lrv ; Freidel:2021ytz ; Himwich:2021dau . However, the infinite number of higher spin currents we get from the loop level soft photon theorems and the symmetries they generate are robust against quantum corrections. It is to been how one can exploit this feature to constrain scattering amplitudes or relate the Celestial holography to string theory in asymptotically flat space-time [^2] We have more to say on this in an upcoming paper..
The paper is organized as follows. In sections II and III we discuss the loop level soft theorems and their representation in the conformal basis. In section-IV we introduce the dipole current in order to write down the one-loop exact soft photon theorem as a Ward-identity. In the following section till section-VII we discuss conformal transformation property of the dipole operator and correlation functions with multiple insertions of dipole operators. In section- VIII we discuss soft photon theorems beyond one loop and show how they give rise to higher spin currents. In section-IX we discuss the algebra of the higher spin currents and point out the role of $k$ in determining the algebra. In section-X we discuss the issue of the closure of the algebra for finite value of $k$ and the possibility of the algebra getting enhanced to the wedge subalgebra of $W_{1+\infty}$. In section-XI we discuss potential problems which arise in determining the value of $k$.
II All loop soft theorems in momentum space
@@
Consider the scattering of massless charged particles in a generic theory of quantum electrodynamics in four spacetime dimensions, where, apart from electromagnetism, no other long-range interaction is present. The soft photon theorem valid at all order in perturbation theory relating the scattering amplitude $\mathcal{A}^{(\text{em})}_{N+1}$, involving $N$ number of finite-energy charged particles with charges $\{e_{a}\}$ and momenta $\{p_{a}\}$, and one outgoing soft photon with polarization $\varepsilon^{h}_{\mu}$, helicity $h$, and momentum $k^{\mu}=-\omega\mathbf{n}^{\mu}$ (where $\omega$ is the photon energy and $\mathbf{n}^{\mu}$ is the unit null vector specifying the emission direction), to the scattering amplitude $\mathcal{A}^{(\text{em})}_{N}$ without the external photon reads: [^3] As is well known, in four spacetime dimensions both $\mathcal{A}^{(\text{em})}_{N+1}$ and $\mathcal{A}^{(\text{em})}_{N}$ are infrared (IR) divergent. However, since the soft factor appears as a ratio of these amplitudes, it is IR finite in the sense that the soft theorem in (1) relates the IR-finite hard parts of the two amplitudes, following the arguments of Sahoo:2018lxl ; Arkani-Hamed:2020gyp ; Krishna:2023fxg .
@@
The soft factor coefficient $\mathbb{S}_{\text{em}}^{(n)}$ has been argued to be $(n+1)$-loop exact, and is derived in Gell-Mann:1954wra ; Low:1954kd ; Weinberg:1965nx for n=1n=-1, Sahoo:2018lxl for $n=0$ and Sahoo:2020ryf for $n=1$ for scattering of massive charged particles and proved to be universal in Krishna:2023fxg . Below, we present their expressions for the scattering of massless charged particles, together with a conjectured form valid for generic $n$, motivated by the analysis in appendix A and inspired by refs. Sahoo:2020ryf ; Karan:2025ndk .
@@
for $n\geq 0$. The soft factors $\mathbb{S}^{(n)}_{\text{em}}$ for $n\geq 0$ presented above are intrinsically quantum contributions to the soft photon theorems, since the classical soft photon factors vanish in the massless limit. [^5] The classical contribution to $\mathbb{S}^{(n)}_{\text{em}}$ for the scattering of massive charged particles was derived in Karan:2025ndk . It vanishes in the massless limit, which is physically expected since, even in the presence of a long-range electromagnetic force, massless charged particles always move at the speed of light in flat spacetime, do not accelerate, and therefore do not radiate. The terminology “classical” and “quantum” follows from Sahoo:2018lxl . In the above expressions, we adopt the mostly positive convention for the Minkowski metric. The charges and energies of incoming particles are taken with their physical signs, while those of outgoing particles acquire an additional negative sign, ensuring that charge and momentum conservation are expressed as $\sum\limits_{a}e_{a}=0$ and $\sum\limits_{a}p_{a}=0$, respectively. From the soft expansion (1) and (2), the soft factor coefficient $\mathbb{S}_{\text{em}}^{(n)}$ can be extracted by using the following projection operator
@@
In the presence of additional long-range interactions mediated by other massless fields, the soft factor coefficients $\mathbb{S}_{\text{em}}^{(n)}$ receive further universal corrections for all $n\geq 0$. For instance, the correction to $\mathbb{S}_{\text{em}}^{(0)}$ due to gravitational interactions was analyzed in Sahoo:2018lxl . In the context of the results presented above, even if we restrict attention to theories in which the only massless fields are the photon and the charged fields appearing as external hard particles, loop diagrams involving the same massless charged particles, specifically ladder and crossed-ladder topologies, can still generate additional contributions to $\mathbb{S}_{\text{em}}^{(n)}$ when three-point interaction vertices among the massless charged particles are allowed. This effect is discussed in a footnote of Appendix A for the case of scalar charged particle scattering. Consequently, the universality of the soft factors $\mathbb{S}_{\text{em}}^{(n)}$ for $n\geq 0$ is more restrictive than that of the leading soft factor $\mathbb{S}_{\text{em}}^{(-1)}$, as discussed in Krishna:2023fxg .

### v3

@@ Abstract
Soft factorization theorems can be reinterpreted as Ward identities for (asymptotic) symmetries of scattering amplitudes in asymptotically flat space-time. In this paper we study the symmetries implied by the all loop soft photon theorems when all the charged particles are highly energetic and the relation $\omega<<m<<E$ holds where $E$ is the typical energy of a charged particle, $m$ is the typical mass and $\omega$ is the soft photon energy. Loop level soft theorems are qualitatively different from the tree level soft theorems because loop level soft factors contain multi-particle sums. If we want to interpret them as Ward identities or define celestial OPE between between soft and hard operators then we need to introduce additional fields which live on the celestial sphere but do not appear as asymptotic states in any scattering experiment. For example, if we want to interpret the one-loop exact $\mathcal{O}(\ln\omega)$ soft theorem for a positive helicity soft photon (with energy $\omega$) as a Ward identity then we need to introduce a pair of antiholomorphic currents on the celestial sphere which transform as a doublet under the $SL(2,\mathbb{R})_{R}$. We call them dipole currents because the corresponding charges measure the monopole and the dipole moment of an electrically charged particle on the celestial sphere. More generally, the soft photon theorem at $\mathcal{O}(\omega^{2j-1}(\ln\omega)^{2j})$ for every $j\in\frac{1}{2}\mathbb{Z}_{+}$ gives rise to $(2j+1)$ antiholomorphic currents which transform in the spin-$j$ representation of the $SL(2,\mathbb{R})_{R}$. These currents exist in the quantum theory because they follow from loop level soft theorems. We argue that under certain circumstances the (classical) algebra of the higher spin currents is the wedge subalgebra of the $w_{1+\infty}$.
@@ I Introduction
In this paper, we study the symmetry interpretation of the soft photon theorems that remain valid even at loop level. Consequently, the symmetries implied by these soft theorems are suppose to be symmetries of the holographic dual celestial CFT2, persisting even after quantum corrections are taken into account. It is well known that the leading soft photon theorem, derived in Gell-Mann:1954wra ; Low:1954kd ; Weinberg:1965nx , is universal and exact at tree level, receiving no loop corrections [^1] Up to a subtlety recently pointed out in Ma:2023gir , arising from a special class of three-loop diagrams in the presence of massless charged particles.. However, since four-dimensional $S$-matrices are infrared divergent due to the long-range interactions mediated by massless particles, deriving finite subleading corrections has long remained a challenge Cachazo:2014dia ; Bern:2014oka ; He:2014bga . By interpreting soft factors as ratios of $S$-matrices with and without an external soft photon, and by systematizing the exponentiation of infrared divergences, it is possible to extract well-defined subleading soft theorems from the IR-finite parts of the ratios of $S$-matrices. This program was successfully carried out in Sahoo:2018lxl , with a proof of universality subsequently established in Krishna:2023fxg . The resulting subleading soft photon theorem appears at $\mathcal{O}(\ln\omega)$ in the soft photon energy $\omega$ and is one loop exact. The $\mathcal{O}(\ln\omega)$ soft factor consists of two distinct contributions, termed the “classical” and “quantum” parts in Sahoo:2018lxl . The classical contribution encodes information about the classical electromagnetic waveform in the low-frequency expansion, as demonstrated through independent classical analyses in Laddha:2018myi ; Saha:2019tub . By contrast, a clear physical interpretation of the quantum contribution remains elusive to date. In this paper we aim to understand it as a Ward identity in the framework of celestial CFT and to uncover the associated symmetry structure.
@@
From a four-dimensional bulk perspective, the classical part of the subleading soft photon theorem at $\mathcal{O}(\ln\omega)$ has already been understood as a Ward identity associated with a large gauge symmetry. [^2] The Celestial OPE of gluons and gravitons at one-loop were studied in Bhardwaj:2022anh ; Krishna:2023ukw . This symmetry acts on charged matter fields as a local enhancement of global phase rotations, referred to as “superphaserotations” in Campiglia:2019wxe ; AtulBhatkar:2019vcb ; Choi:2025mzg . In the present work we provide an interpretation of the quantum contribution to the $\mathcal{O}(\ln\omega)$ soft photon theorem in QED as a symmetry of the dual Celestial CFT2 in the high energy limit. Furthermore, we analyze higher subleading corrections to the soft photon theorem at orders $\omega^{n}(\ln\omega)^{n+1}$ as conjectured and explicitly derived up to sub-subleading order in Sahoo:2020ryf ; Karan:2025ndk , and argued to be $(n+1)$-loop exact. In the high energy limit the classical part of the soft theorem is subleading to the quantum part.
@@
Loop level soft theorems are qualitatively different from the tree level ones. It turns out that if we want to interpret them as Ward identities of symmetries in a Celestial CFT2 or define a Celestial OPE between a hard and soft operators then we need to introduce additional fields on the celestial sphere which do not appear as asymptotic states in scattering experiments. In particular, they are not related to the soft gauge bosons in any obvious manner. Now since there are an infinite number of loop level soft photon theorems we get an infinite number of chiral currents in this way. For every $j\in\frac{1}{2}\mathbb{Z_{+}}$ the soft theorem at $\mathcal{O}(\omega^{2j-1}(\ln\omega)^{2j})$ gives rise to $(2j+1)$ chiral currents which transform in the finite dimensional spin-$j$ representation of the $SL(2,\mathbb{R})_{R}$. The charges corresponding to the $j=1/2$ currents measure the monopole and the dipole moment of an electrically charged particle on the celestial sphere and hence called a dipole current. The other spin-$j$ currents can be constructed by taking the (normal ordered) product of $2j$ dipole currents. Depending on the value of a real parameter $k$ which appears in the OPE of the dipole currents, the algebra of the higher spin currents can be Abelian ($k=0$) or the wedge subalgebra ($k<<1$) of the $w_{1+\infty}$.
@@
Appearance of higher spin currents is not new in the context of Celestial holography. Both soft gluons and gravitons generate infinite dimensional higher spin algebras known as the $S$-algebra and the wedge subalgebra of $w_{1+\infty}$ Strominger:2021mtt ; Adamo:2021lrv ; Freidel:2021ytz ; Himwich:2021dau . However, the infinite number of higher spin currents we get from the loop level soft photon theorems and the symmetries they generate are robust against quantum corrections, at least in the high energy limit. It remains to be seen how one can exploit this feature to constrain scattering amplitudes or relate celestial holography to string theory in asymptotically flat space-time. [^3] We have more to say on this in an upcoming paper.
Note Added: In a previous version of the paper instead of the high energy limit we talked about the massless limit of the QED. Although in many cases these two limits are synonymous, in the present context this is not so. There are two reasons. Firstly, there are two low energy scales in the problem. One is the mass $m$ of the charged particle and the other one is the energy $\omega$ of the soft photon. The results of Sahoo:2018lxl were derived under the assumption that $\omega<<m$ and so massless limit cannot be taken. In other words, $\omega$ is the smallest energy scale in the problem and this should be strictly maintained at every stage of any limiting procedure. The second, and more fundamental, reason is that in the massless QED the physical charge goes to zero in the infrared and therefore soft symmetries cannot be defined. We are grateful to Andrew Strominger for pointing this out to us.
The high energy limit that is discussed in the following section does not suffer from any of these problems and all the technical results presented in the previous version remain unchanged.
II All loop soft theorems and high energy limit
@@
Consider the scattering of massive charged particles in a generic theory of quantum electrodynamics in four spacetime dimensions, where, apart from electromagnetism, no other long-range interaction is present. The soft photon theorem valid at all order in perturbation theory relating the scattering amplitude $\mathcal{A}^{(\text{em})}_{N+1}$, involving $N$ number of finite-energy charged particles with charges $\{e_{a}\}$ and momenta $\{p_{a}\}$, and one outgoing soft photon with polarization $\varepsilon^{h}_{\mu}$, helicity $h$, and momentum $k^{\mu}=-\omega\mathbf{n}^{\mu}$ (where $\omega$ is the photon energy and $\mathbf{n}^{\mu}$ is the unit null vector specifying the emission direction), to the scattering amplitude $\mathcal{A}^{(\text{em})}_{N}$ without the external photon reads: [^4] As is well known, in four spacetime dimensions both $\mathcal{A}^{(\text{em})}_{N+1}$ and $\mathcal{A}^{(\text{em})}_{N}$ are infrared (IR) divergent. However, since the soft factor appears as a ratio of these amplitudes, it is IR finite in the sense that the soft theorem in (1) relates the IR-finite hard parts of the two amplitudes, following the arguments of Sahoo:2018lxl ; Arkani-Hamed:2020gyp ; Krishna:2023fxg ., [^5] In this paper, we adopt the mostly positive convention for the Minkowski metric. The charges and energies of incoming particles are taken with their physical signs, while those of outgoing particles acquire an additional negative sign, ensuring that charge and momentum conservation are expressed as $\sum\limits_{a}e_{a}=0$ and $\sum\limits_{a}p_{a}=0$, respectively.
@@
The soft factor $\mathbb{S}_{\text{em}}^{(n)}$ has been argued to be $(n+1)$-loop exact. The soft factor $\mathbb{S}_{\text{em}}^{(-1)}$ was derived in Gell-Mann:1954wra ; Low:1954kd ; Weinberg:1965nx , $\mathbb{S}_{\text{em}}^{(0)}$ in Sahoo:2018lxl , and $\mathbb{S}_{\text{em}}^{(1)}$ in Sahoo:2020ryf , for the scattering of massive charged particles. The universality of $\mathbb{S}_{\text{em}}^{(0)}$ was established in Krishna:2023fxg .
II.1 High energy limit
In this paper we want to study the soft expansion (2) in the high energy limit
$\boxed{\omega^{2}<<m^{2}<<|p_{a}.p_{b}|\sim E^{2}},\quad\forall a\neq b,$ (3)
where $m$ represents the typical mass of a charged particle, $E$ is the characteristic energy scale of the hard charged particles and $\omega$ is the energy of the soft photon.
In the high energy limit (3) the expressions for the soft factors simplify considerably. For example, the $\mathcal{O}(\ln\omega)$ soft factor given by Sahoo:2018lxl
$\begin{gathered}\mathbb{S}^{(0)}_{\text{em}}=\frac{i}{4\pi}\sum_{a=1}^{N}\sum^{N}_{\begin{subarray}{c}b=1\\ b\neq a,\eta_{a}\eta_{b}=1\end{subarray}}e_{a}^{2}e_{b}\frac{\epsilon_{\mu}\mathbf{n}_{\rho}}{p_{a}.\mathbf{n}}\frac{m_{a}^{2}m_{b}^{2}\left[p^{\mu}_{a}p^{\rho}_{b}-p^{\mu}_{b}p^{\rho}_{a}\right]}{\left[(p_{a}.p_{b})^{2}-m_{a}^{2}m_{b}^{2}\right]^{\frac{3}{2}}}\\ +\frac{1}{8\pi^{2}}\sum^{N}_{\begin{subarray}{c}a,b=1\\ a\neq b\end{subarray}}e_{a}^{2}e_{b}\log\left[\frac{p_{a}.p_{b}+\sqrt{(p_{a}.p_{b})^{2}-m_{a}^{2}m_{b}^{2}}}{p_{a}.p_{b}-\sqrt{(p_{a}.p_{b})^{2}-m_{a}^{2}m_{b}^{2}}}\right]\frac{m_{a}^{2}m_{b}^{2}}{\left((p_{a}.p_{b})^{2}-m_{a}^{2}m_{b}^{2}\right)^{3/2}}\left(-\epsilon.p_{b}+\frac{\epsilon.p_{a}}{\mathbf{n}.p_{a}}\mathbf{n}.p_{b}\right)\\ -\frac{1}{4\pi^{2}}\sum^{N}_{\begin{subarray}{c}a,b=1\\ a\neq b\end{subarray}}e_{a}^{2}e_{b}\frac{p_{a}.p_{b}}{(p_{a}.p_{b})^{2}-m_{a}^{2}m_{b}^{2}}\left(-\epsilon.p_{b}+\frac{\epsilon.p_{a}}{\mathbf{n}.p_{a}}\mathbf{n}.p_{b}\right)\end{gathered}$ (4)
can be expanded in the limit (3) and the result is given by
$\begin{gathered}\mathbb{S}^{(0)}_{\text{em}}=-\frac{1}{4\pi^{2}}\sum_{\begin{subarray}{c}a,b=1\\ b\neq a\end{subarray}}^{N}\frac{e^{2}_{a}e_{b}}{p_{a}.p_{b}}\frac{\varepsilon_{\mu}\mathbf{n}_{\nu}}{p_{a}.\mathbf{n}}\left(p_{a}^{\mu}p_{b}^{\nu}-p_{a}^{\nu}p_{b}^{\mu}\right)+\mathcal{O}\left(\frac{m^{4}}{E^{5}}\log\frac{m^{4}}{E^{4}},\frac{m^{4}}{E^{5}}\right).\end{gathered}$ (5)
We observe that the soft factor $\mathbb{S}^{(0)}_{\text{em}}$ contains a leading contribution of order $\mathcal{O}(1/E)$ which is a rational function of the hard particle momenta. In addition, it includes power-law and logarithmic correction terms, as indicated inside the parentheses of the $\mathcal{O}(\cdots)$ symbol. These terms are subleading in the limit defined in (3). Throughout this paper, we restrict our attention to the leading rational contribution appearing in (5).
Now in QED the high energy limit is not really well defined because of the existence of the Landau pole. We assume that the energies of the hard particles are well below that but they are much greater than the typical mass of a charged particle so that we can neglect the logarithmic and power law suppressed terms in (5) to a good approximation. However, it should be kept in mind that there is no region of parameter space where the correction terms are zero. It will be very interesting to study the soft symmetry interpretation of the subleading terms in the high energy expansion. We leave it for future investigation.
Now motivated by the analysis in appendix A and inspired by refs. Sahoo:2020ryf ; Karan:2025ndk we can write down the expressions for the leading terms of the logarithmic soft theorems in the high energy limit (3) as
@@ II.1 High energy limit
for $n\geq 0$. The $\mathbb{S}^{(n)}_{\text{em}}$ for $n\geq 0$ presented above are intrinsically quantum contributions to the soft photon theorems since the classical contributions are subleading in the high energy limit. [^7] The classical contribution to $\mathbb{S}^{(n)}_{\text{em}}$ for the scattering of massive charged particles was derived in Karan:2025ndk . It vanishes in the massless limit, which is physically expected since, even in the presence of a long-range electromagnetic force, massless charged particles always move at the speed of light in flat spacetime, do not accelerate, and therefore do not radiate. The terminology “classical” and “quantum” follows from Sahoo:2018lxl . From the soft expansion (1) and (2), the soft factor coefficient $\mathbb{S}_{\text{em}}^{(n)}$ can be extracted by using the following projection operator